Rob Colley set a record in "pole-sitting" when he spent 42 days in a barrel at the top of a flagpole with a height of 43m

8.
Rob Colley set a record in "pole-sitting" when he spent 42 days in a barrel at the top of a flagpole with a height of 43m. Supposed a friend wanting to deliver an ice-cream sandwich to Colley throws the ice cream straight up woth just enough speed to reach the barrel. How long does it take the ice-cream sandwich to reach the barrel?
6.
The Westin Stamford Hotel in Detroit is 228m. tall. If a worker on the roof drops a sandwich, how long does it take the sandwich to hit the ground, assuming there is no air resistance? How would air resistance affect the answer?
4.
The tallest roller coaster in the world is Desperado in Nevada. It has a lift height of 64m. If an archer shoots an arrow straight up in the air and the arrow passes the top of the roller coaster 3.0seconds the arrow is shot, what is the initial speed of the arrow?
2.
Brian Berg of Iowa built a house of cards 4.88 meters tall. Suppose Berg throws a ball from ground level with a velocity of 9.98 m/s straight up. What is the velocity of the ball as it firsts passes the top of the card house?
Answer #1

6 ans:

d=a*t^2/2


-228=-4.9*t^2

t^2=46.53......

t=6.82 seconds


Air resistance would put an upward force on the sandwich, causing it to accelerate downward at a slower rate, so the time would be longer.


answered by: nobody
Answer #2

4 ans: 

Time (t) = 3s

Displacement (s) = 64m

Acceleration (g) = -9.8m/s^2

Initial speed (u) = ?

Use the following equation:


s = ut + 1/2 at^2

64 = u*3 + 0.5 * -9.8 * 3^2

64 = 3u - 44.1

u = 36.03m/s


answered by: nobody